OpenStudy (anonymous):

Get a medal :3 Evaluate the expression when p = -16 and q = 4 4p/q A. –16 B. –3 C. 3 D. 16

OpenStudy (anonymous):

\[\frac{ 4p }{ q }\]

OpenStudy (anonymous):

The answer is A. Explanation: -16*4=-64 -64/4=-16 Answer: -16

OpenStudy (anonymous):

A.